### v1

# NarrativeBridge: Enhancing Video Captioning with Causal-Temporal Narrative

Abstract
Existing video captioning benchmarks and models lack coherent representations of causal-temporal narrative, which is sequences of events linked through cause and effect, unfolding over time and driven by characters or agents. This lack of narrative restricts models’ ability to generate text descriptions that capture the causal and temporal dynamics inherent in video content. To address this gap, we propose NarrativeBridge, an approach comprising of: (1) a novel Causal-Temporal Narrative (CTN) captions benchmark generated using a large language model and few-shot prompting, explicitly encoding cause-effect temporal relationships in video descriptions, evaluated automatically to ensure caption quality and relevance; and (2) a dedicated Cause-Effect Network (CEN) architecture with separate encoders for capturing cause and effect dynamics independently, enabling effective learning and generation of captions with causal-temporal narrative. Extensive experiments demonstrate that CEN is more accurate in articulating the causal and temporal aspects of video content than the second best model (GIT): 17.88 and 17.44 CIDEr on the MSVD and MSR-VTT datasets, respectively. The proposed framework understands and generates nuanced text descriptions with intricate causal-temporal narrative structures present in videos, addressing a critical limitation in video captioning. For project details, visit https://narrativebridge.github.io/.
1 Introduction
Video captioning aims to generate textual descriptions that capture the visual information and temporal dynamics in videos [48]. Research has primarily focused on developing new methods to improve the accuracy of video captioning models on well-established benchmark datasets, MSR-VTT [55] and MSVD [7]. State-of-the-art (SOTA) approaches [19, 44, 56, 32, 51, 9] proposed new architectures to better align the generated captions with the provided ground-truth. However, these efforts enhance models’ accuracy on existing evaluation criteria, without modifying the benchmark datasets or their ground truth annotations to address the lack of coherent representations of causal-temporal narrative - construction and interpretation of a sequence of events linked through cause and effect, unfolding over time and space, and often driven by entities (characters or agents) acting with intention [52].
As discussed by Wilkens et al. [52], causal relationships [16] are scenarios where one event (cause) directly influences the occurrence of another event (effect), while temporal understanding [4] involves recognizing the chronological order of events, a crucial component in establishing causality. Existing ground-truth captions in popular benchmarks lack causal, temporal and narrative information. Figure 1 illustrates the importance of causal-temporal narrative on the MSR-VTT dataset [55]. In the input video sequence, a car is first shown driving recklessly through an open field and flipping over, which represents the cause event. This is then followed by a scene showing the severely damaged car, and subsequently, a group of guys starting to play beer pong, which represents the effect event. The causal-temporal narrative would connect these events in a temporal order, linking the reckless driving and resulting car crash to the group’s subsequent behavior of playing beer pong. This example aligns with the definition of narrative provided by Wilkens et al. [52], which emphasizes the role of cause-effect relationships, the perception of narrativity in videos, and highlights the temporal sequence of events that is crucial for understanding the causal-temporal narrative. Figure 1 highlights the limitations of the original captions in existing benchmark datasets, such as MSR-VTT [55]. The original captions focus on isolated events or actions, such as "a car flipping over" or "a monster truck flips on its side," lacking contextual narrative and causal-temporal relationships between events. Consequently, models trained on these captions suffer from the same limitations.
Figure 1: Comparison of Original captions vs. Causal-Temporal Narrative (CTN) caption to illustrate the inclusion of causal-temporal narrative.
To bridge this gap, we introduce NarrativeBridge, a novel framework encompassing a new benchmark and architecture tailored for causal-temporal narrative learning in video captioning. Our Causal-Temporal Narrative (CTN) captions benchmark leverages a large language model (LLM) and few-shot prompting to generate enhanced video descriptions that explicitly encode cause-effect relationships and temporal sequences, as exemplified in Figure 1. Our CTN captions benchmark, generated using the approach detailed in Section 3, establishes a clear connection between the cause (reckless driving) and the effect (damaged car and subsequent behavior of the group). By generating a single, coherent caption that encapsulates the causal-temporal narrative of the video, our CTN captions benchmark enables models to better understand and articulate the causality, sequence, and significance of events within the broader video context [52]. This approach addresses the limitations of existing benchmark datasets and emphasizes the importance of incorporating causal-temporal narrative understanding into video captioning models to generate accurate, informative, and contextually relevant descriptions. To ensure the quality and relevance of the generated captions, we employ an automatic evaluation framework that compares the CTN captions with the video content, keeping or discarding them based on a score threshold. This approach addresses the limitations of existing benchmark datasets and emphasizes the importance of incorporating causal-temporal narrative understanding into video captioning models to generate accurate, informative, and contextually relevant descriptions.
Existing SOTA video captioning methods that use different architectures such as LSTM [32], GNN [17], and Transformer [51], struggle to effectively learn causal-temporal narrative from the CTN captions. These architectures are designed to capture the overall semantics in videos but lack dedicated mechanisms to explicitly model the cause-effect relationships and temporal sequences. As a result, the captions generated by these methods fail to articulate the complex causal-temporal dynamics in the videos, as demonstrated in results (Section 4.4). To address this challenge, we propose the Cause-Effect Network (CEN), an architecture specifically designed to capture cause and effect dynamics using dedicated encoders. By separately encoding the cause and effect components of the video and then combining them to generate the final caption, CEN is able to better understand and articulate the causal-temporal narrative. The primary contributions of our work are:
  • Introducing for the first time CTN captions benchmark with high-quality, coherent causal-temporal narrative captions, automatically evaluated, to address the lack of coherent representations in video captioning benchmarks.
  • A novel CEN network for learning causal-temporal narrative from videos to address limitation of SOTA video captioning methods to learn narrative information on CTN.
  • Extensive experiments demonstrating CEN’s superior performance on CTN, setting a new SOTA in causal-temporal narrative learning.
Our work marks a significant step forward in video captioning research by explicitly addressing the challenges of causal-temporal narrative understanding. The CTN captions benchmark and the CEN architecture provide a comprehensive framework for learning and generating contextually relevant and temporally coherent video descriptions. This framework has the potential to drive the development of more sophisticated video understanding systems and enhance various applications, such as video summarization, content retrieval, and human-robot interaction.
2 Related Work
2.1 Benchmarks
MSVD [7] is a benchmark focused on human activities, providing a platform for evaluating video captioning models. The captions in MSVD often describe the observable actions without delving into the underlying motivations or the cause-effect relationships between the events. MSR-VTT [55] is a large-scale benchmark with diverse video content, encompassing a wide range of topics and genres. The captions often focus on describing the observable content without capturing the causal links between the events or the temporal progression of the narrative. As a result, models trained on MSVD and MSR-VTT may struggle to generate descriptions that accurately reflect the causal and temporal dynamics in the videos.
While recent benchmarks like NExT-QA [54] and EgoSchema [30] have made strides in incorporating causal and temporal reasoning, they focus on specific aspects of video understanding, such as question-answering and long-form video comprehension. NExT-QA [54] introduces multi-choice and open-ended question-answering tasks that target causal and temporal action reasoning but do not address the generation of holistic video descriptions. EgoSchema [30], on the other hand, emphasizes long-form video understanding and temporal reasoning but does not explicitly focus on causal-temporal narrative for video captioning.
Similarly, efforts like VCR [60], V2C [12], and Motivation [49] integrate causality into their analysis of visual description or question-answering, relying heavily on commonsense reasoning for generating predictions. VCR [60] focuses on visual commonsense reasoning, V2C [12] aims to generate commonsense descriptions for video captioning, and Motivation [49] explores the prediction of motivations behind actions in videos. However, these works primarily rely on commonsense reasoning and do not delve into the causal and temporal structures underpinning video narratives. We introduce a new benchmark CTN specifically designed to capture and evaluate causal-temporal narrative in video captioning.
2.2 Video Captioning
Video captioning techniques have evolved from LSTM-based [13, 42, 32] frameworks to the latest designs using SOTA GNNs [17, 62, 34] and Transformers [51, 24, 58], with a focus on enhancing the complexity of captions through the injection of multimodal data. Despite these advancements, current architectures often struggle to capture the intricate temporal sequences and causal relationships in video storytelling. To bridge this gap, video captioning can benefit from cross-fertilization with ideas and strategies developed in related fields, such as action recognition [43, 50, 21, 53, 8, 14, 35, 38, 2, 36, 61], event localization [45, 25, 11, 26], and question-answering [1, 18, 39, 15, 59, 22, 40, 33]. The integration of causal reasoning [27, 57] has shown promise in enhancing the ability of neural networks to discern causal relationships, leading to improved performance in image captioning [27] and visual question answering [57]. However, current SOTA models still struggle to effectively handle the narrative complexity in videos, particularly in terms of causal and temporal progression. To address this limitation, there is a need to develop new video captioning models that can overcome the shortcomings of current frameworks.
In light of these challenges, our work aims to bridge the gap in existing video captioning benchmarks and models by introducing NarrativeBridge, a comprehensive framework that encompasses the CTN captions benchmark and the CEN architecture. By explicitly addressing the limitations of current approaches and providing a platform for causal-temporal narrative learning, our work paves the way for the development of more sophisticated video captioning models capable of generating contextually relevant and temporally coherent descriptions.
Figure 2: CTN caption generation pipeline. $\theta$ indicates a threshold.
3 Method
NarrativeBridge is a comprehensive approach to video captioning that addresses the limitations of existing benchmarks and models in capturing causal-temporal narrative. It consists of two key interconnected components: (i) the CTN captions benchmark, which provides a rich representation of cause-and-effect relationships and event sequences in video content, and (ii) the CEN architecture, which is specifically designed to learn and articulate these causal-temporal narrative elements.
Problem Statement: Existing video captioning benchmarks and models struggle to coherently capture the intricate causal-temporal narrative inherent in video content. As shown in Figure 1, original captions from benchmark datasets often focus on isolated events, failing to capture the causal and temporal relationships that drive the video’s narrative. SOTA video captioning methods also fall short in articulating the cause-and-effect relationships and temporal sequences that drive the events in the video (see Section 4.4).
You are an advanced language model tasked with generating causal-temporal narrative captions for a
video. However, you cannot directly access the video itself. Instead, you will be provided with a
series of captions that outline the key events and scenes in the video. Your task is to generate a
concise Cause and Effect scenario, based on the information provided in the descriptive captions.
Be careful, your generated Cause and Effect statements should fulfill the following requirements:
1. Your narrative should be grounded in the information provided by the descriptive captions.
2. Cause and Effect scenario is relevant.
3. It should not introduce any new events or details not mentioned.
4. Avoid implying conclusions.
5. Maintain temporal consistency with the provided captions.
6. Use plain English and direct sentences.
7. Cause and Effect statements each limited to a maximum of 15 words.
8. Do not include any additional text before or after the JSON object.
Here are the examples of Cause and Effect:
[Examples]:
[{’Cause’: the student overslept due to a malfunctioning alarm clock’, Effect’: missed catching the
bus to school’}, {’Cause’: she absentmindedly skipped applying moisturizer after taking a long hot
shower’, Effect’: her skin became dry and flaky’}, {’Cause’: he carelessly neglected taking his
prescribed allergy medication’, Effect’: suffered a severe sneezing fit’}, {’Cause’: the exhausted
soccer player recklessly fouled an opponent in the penalty area’, Effect’: the opposing team was
awarded a crucial penalty kick’}, {’Cause’: due to unforeseen road closures they found themselves
stuck in heavy traffic’’, Effect’: missed out on experiencing the opening act of the concert’}]
Now please generate only one Cause and Effect presented in a JSON format based on the following
descriptive captions.
[Descriptive Captions]:
<descriptive_captions>
[Causal Temporal Narrative]:
Prompt 1: LLM Prompt used in our CTN captions generation.
3.1 CTN Captions Benchmark
To address the challenges of existing benchmarks in representing causal and temporal relationships within video content, we propose an approach that harnesses the potential of LLMs through the few-shot prompting technique [5]. Our method generates CTN captions without the need for model fine-tuning, leveraging the inherent generative capabilities of LLMs to produce captions that encapsulate causal-temporal narrative structures. Figure 2 shows the CTN caption generation pipeline, which consists of two key steps: prompt design, LLM-based caption generation and evaluation.
Prompt Design: The prompt design step is crucial in guiding the LLM to generate causal-temporal narrative captions. The input to the CTN caption generation are the original video captions from existing benchmarks. We design a prompt that include a small set of carefully selected example captions, illustrating the desired output structure and highlighting the key aspects of causal-temporal narrative. We design Prompt 1 which guides LLM to generate accurate and concise causal-temporal narrative. The prompt sets clear requirements for grounding the captions in the provided descriptive context, maintains temporal consistency, and avoids unsupported details or conclusions. The illustrative examples demonstrate the desired format, facilitating the generation of plainly written, length-constrained Cause and Effect statements that effectively capture the video’s causal-temporal narrative. By explicitly outlining instructions and constraints, the carefully designed prompt steers the LLM’s generative capabilities towards producing contextually relevant and narratively coherent CTN captions. Further details are provided in Appendix B and C.
CTN caption Generation and Evaluation: We send Prompt 1 into the Mixtral of Experts LLM [20], which generates CTN captions that accurately reflect the causal-temporal narrative in the video. The LLM’s advanced natural language understanding and generation capabilities, combined with the few-shot prompting approach, enable the production of captions that encapsulate complex causal and temporal relationships. To ensure the quality and relevance of the generated CTN captions, we employ the EMScore [41] metric for evaluation. EMScore directly measures the consistency between a caption and the video content, and has been shown to be more effective than other metrics [10] in evaluating video relevance without referenced captions. We set a threshold ($\theta$ = 0.2) for the EMScore value, indicating adequate relevance to the video. Captions with an EMScore above the $\theta$ are kept, while those below the $\theta$ are discarded, and the LLM generates a new caption. This iterative refinement process continues until a caption meets the EMScore threshold, ensuring that the final captions efficiently describe the relevant events in the video. Further details are provided in Appendix D.
Figure 3: The two-stage Cause-Effect Network (CEN) architecture. Stage 1: Separate Cause ($E_{cause}$) and Effect ($E_{effect}$) video encoders, pretrained using CLIP-ViT, learn specialized video representations. Corresponding text encoders ($T_{cause}$ and $T_{effect}$) encode the cause and effect portions of the CTN caption. Contrastive losses are applied to align the video and text embeddings. Stage 2: The learned cause and effect video features are encoded separately ($Enc_{cause}$ and $Enc_{effect}$) and concatenated before being input to the decoder, which generates the final CTN caption.
3.2 Proposed Cause-Effect Network (CEN)
Existing SOTA video captioning methods based on LSTMs[32], GNNs[17], and Transformers[51] lack dedicated mechanisms to explicitly model the intricate causal-temporal narrative sequences. Our proposed CEN employs a two-stage approach to address this limitation. In Stage 1 (Section 3.2.1), CEN learns specialized cause and effect video representations using separate encoders trained on the corresponding portions of the CTN captions. The input is raw video frames, and the output is the learned cause and effect features extracted from the encoders. The learned video features from Stage 1 are then passed to Stage 2 (Section 3.2.2), which utilizes a sequence-to-sequence transformer-based network with two new encoders to generate the final captions. The parameters of stage 1 encoders are frozen at this stage. The cause and effect video features are separately encoded and then concatenated before being input to the decoder. This allows the model to synthesize the specialized representations learned in Stage 1 and generate captions that accurately capture the causal and temporal relationships in the videos.
3.2.1 Stage 1: Causal Visual Encoding
For the CEN network to learn causal-temporal narrative, we design two separate visual encoding models trained on two parts of the captions (see Figure 3).
(1) Cause Video Encoder $E_{cause}$: Trained on the cause captions part describing the initiating events;
(2) Effect Video Encoder $E_{effect}$: Trained on the effect part of the captions capturing the consequential outcomes.
We instantiate $E_{cause}$ and $E_{effect}$ as two separate instances of the CLIP-ViT (ViT-B/32) model [37], adapted from CLIP4Clip [29]. For both Cause and Effect text encoders ($T_{cause}$ and $T_{effect}$ respectively), we employ a 12-layer transformer with a hidden size of $512$ and $8$ attention heads, using the weights from the pre-trained CLIP [37] text encoder. As in CLIP [37], the [EOS] token’s representation from the last layer of the text encoders is used as the feature representation for the input text. Similar to CLIP4Clip, we adopt cosine similarity to measure the similarity between the video embedding (mean pooled) and the text embedding. Given a video embedding $\hat{r}_{i}$ and a text embedding $t_{j}$, the similarity function is defined as $s(\hat{r}_{i},t_{j})=\frac{t_{j}^{\top}\hat{r}_{i}}{|t_{j}||\hat{r}_{i}|}$. During training, each model is optimized using a contrastive loss, as in CLIP [37], over its respective cause/effect portion from CTN captions. For a batch of $N$ video-cause text pairs $(v_{i},c_{i})$ and video-effect text pairs $(v_{i},e_{i})$, the contrastive losses for the cause and effect video encoders are defined as:
$\mathcal{L}_{cause}=\mathcal{L}_{v2t}(E_{cause}(v_{i}),T_{cause}(c_{i}))+\mathcal{L}_{t2v}(E_{cause}(v_{i}),T_{cause}(c_{i})),$ (1)
$\mathcal{L}_{effect}=\mathcal{L}_{v2t}(E_{effect}(v_{i}),T_{effect}(e_{i}))+\mathcal{L}_{t2v}(E_{effect}(v_{i}),T_{effect}(e_{i})),$ (2)
where $\mathcal{L}_{v2t}$ and $\mathcal{L}_{t2v}$ are the video-to-text and text-to-video contrastive losses, respectively, defined as: $\mathcal{L}_{v2t}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(s(\hat{r}_{i},t_{i}))}{\sum_{j=1}^{N}\exp(s(\hat{r}_{i},t_{j}))},$ (3) $\mathcal{L}_{t2v}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(s(\hat{r}_{i},t_{i}))}{\sum_{j=1}^{N}\exp(s(\hat{r}_{j},t_{i}))}$ (4)
After training, we freeze the weights of $E_{cause}$ and $E_{effect}$ for the second stage, where we extract the cause and effect video features, respectively.
3.2.2 Stage 2: CTN Caption Generation
In the second stage (see Figure 3), we employ a sequence-to-sequence transformer-based network, consisting of two separate encoders ($Enc_{cause}$ and $Enc_{effect}$) for processing cause and effect video features, respectively, and a decoder $Dec$ for generating the final CTN captions. The encoded cause and effect video representations are concatenated before being sent to the decoder. Both encoders and the decoder are initialized with the weights from the pre-trained Uni-VL model [28]. Given a video $v$, we first extract the cause and effect video features using the pre-trained $E_{cause}$ and $E_{effect}$ encoders: $F_{cause}=E_{cause}(v),$ (5) $F_{effect}=E_{effect}(v).$ (6)
The extracted features $F_{cause}$ and $F_{effect}$ are separately encoded using $Enc_{cause}$ and $Enc_{effect}$, each consisting of two transformer layers, to obtain the respective encoded representations: $h_{cause}=Enc_{cause}({F}_{cause})$ and $h_{effect}=Enc_{effect}({F}_{effect})$.
The encoded cause and effect representations are concatenated to form a single representation $h_{concat}=[h_{cause};h_{effect}]$, which is then input to the decoder $Dec$, consisting of two transformer layers, to generate the CTN caption $\hat{y}$ by attending to $h_{concat}$ at each time step $t$:
$\hat{y}_{t}=Dec(h_{concat}],\hat{y}_{<t}),$ (7)
where $\hat{y}_{<t}$ denotes the previously generated words. For training and evaluation, we combine the cause and effect parts of the ground-truth CTN caption with a space in between to form a single, coherent caption. This combined caption serves as the target for the decoder during training and the reference for evaluating the generated captions using diffferent evaluation metrics. The entire network is trained end-to-end using the cross-entropy loss between the generated caption $\hat{y}$ and the ground-truth CTN caption $y$:
$\mathcal{L}_{caption}=-\frac{1}{T}\sum_{t=1}^{T}\sum_{c=1}^{|V|}y_{t,c}\log(\hat{y}_{t,c}),$ (8)
where: $T$ is the length of the caption. $|V|$ is the size of the vocabulary. $y_{t,c}$ is the ground-truth label for the $c$-th word at time step $t$ (1 if the word is present, 0 otherwise). $\hat{y}_{t,c}$ is the predicted probability of the $c$-th word at time step $t$. Finally, our total loss $\mathcal{L}_{total}$ is:
$\mathcal{L}_{total}=\mathcal{L}_{cause}+\mathcal{L}_{effect}+\mathcal{L}_{caption}$ (9)
This two-stage approach first encodes the causal aspects into the video representations, facilitating the generation of coherent causal-temporal narrative descriptions in the second stage (see Experiments section 4) to generate CTN captions as final output.
4 Experiments
We conduct extensive experiments to validate our proposed CEN architecture and the new CTN video captioning benchmark. This section details the datasets used, the experimental setup, and the evaluation metrics. We also present comprehensive results and evaluation against the state-of-the-art video captioning methods demonstrating the performance of our network on the CTN benchmark.
4.1 Datasets
We evaluate our CTN captions and CEN on two widely-used video captioning datasets: 1) MSR-VTT [55], consists of 10,000 videos with 200,000 human-annotated captions, and MSVD [7], with 1,970 videos focused on human activities with 78,800 captions. The captions in these datasets lack explicit representations of the causal relationships, temporal dynamics, motivations, intentions, and the narrative of events and human actions in the videos. To address these limitations, we generate CTN captions using the Mixtral of Experts LLM [20] as detailed in Section 3.1.
4.2 Evaluation Metrics
For quantifying the performance of our method CEN against the SOTA methods, we use three metrics as per [6]; CIDEr (C) [47]: Measures alignment of generated captions with references by highlighting frequently occurring terms, capturing the ability to reproduce salient causal-temporal narrative elements. ROUGE-L (R-L) [23]: Evaluates the longest common subsequence between generated and reference captions, considering both precision and recall for assessing semantic similarity, including temporal dynamics and causal relationships. SPICE (S) [3]: Evaluates semantic quality by considering the overlap of scene graphs between generated and reference captions, effectively assessing the ability to capture and express underlying causal relationships and event sequences.
4.3 Implementation Details
We generate the CTN captions using the Mixtral of Experts LLM [20], running on A100-80GB GPUs. All models, including our proposed CEN architecture and the SOTA models (SEM-POS [32], AKGNN [17], and GIT [51]), are implemented using PyTorch. For the SOTA models, we follow the hyperparameter settings specified in their respective methods for training on the MSR-VTT and MSVD datasets. Our CEN model is trained using the Adam optimizer with learning rates of $1\times 10^{-4}$ (stage 1) and $1\times 10^{-6}$ (stage 2) and a batch size of 64 for 10 epochs (stage 1) and 50 epochs (stage 2). Further details are provided in the Appendix A.
4.4 Results and Discussion
We comprehensively evaluate the performance of our proposed CEN architecture against several SOTA methods, including SEM-POS [32], AKGNN [17], and GIT [51], on our CTN captions generated for MSR-VTT [55] and MSVD [7] benchmark datasets.
Quantitative Results: CEN outperformed all SOTA methods across all metrics and datasets as shown in Table 2. On MSVD, CEN surpassed the next best (GIT) by 3.95 ROUGE-L, 17.88 CIDEr, and 3.67 SPICE points. On MSR-VTT, CEN led GIT by 3.39 ROUGE-L, 17.44 CIDEr, and 2.06 SPICE points. These significant gains highlight CEN’s effectiveness in capturing causal narratives and temporal dynamics. To gain further insights into the effectiveness of the CTN captions in capturing causal-temporal narratives, we visualize the video feature representations learned by CLIP-ViT encoders trained on CTN and Original captions using UMAP [31] dimensionality reduction from high dimension onto a 2D plane. In Figure 4(a), we compare the representations learned from CTN captions (red) and Original captions (blue) on the MSR-VTT dataset. The non-overlapping feature spaces indicate that CTN captions capture the causal and temporal relationships not present in the original captions. In Figure 4(b), we visualize the representations learned from the cause (black) and effect (orange) parts of the CTN captions. The near-complete overlap suggests a strong correlation between the cause and effect components, aligning with the inherent structure of causal-temporal narratives and supporting the design of the CEN architecture.
Table 1: Comparison of our CEN architecture against SOTA methods on the MSVD and MSR-VTT datasets. The best results in each category are in bold. R-L, C, and S denote ROUGE-L, CIDEr, and SPICE scores, respectively. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) SEM-POS [32] 25.39 37.16 14.46 20.11 26.01 12.09 AKGNN [17] 25.11 35.08 14.55 21.42 25.90 11.99 GIT [51] 27.51 45.63 15.58 24.51 32.43 13.70 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76 Table 2: Ablation study results on the MSVD and MSR-VTT datasets. $\text{E}_{\text{JointCE}}$, Only $\text{E}_{\text{cause}}$ and Only $\text{E}_{\text{effect}}$ are variants of our CEN architecture, while Zero Shot v.v. and Fine-tune v.v. represent cross-dataset evaluation settings. The best results in each category are in bold. Method MSVD MSRVTT R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) R-L ()($\uparrow$)( ↑ ) C ()($\uparrow$)( ↑ ) S ()($\uparrow$)( ↑ ) $\text{E}_{\text{JointCE}}$ 30.93 55.72 17.04 27.34 45.97 15.07 Only $\text{E}_{\text{cause}}$ 30.72 56.42 18.43 27.19 47.10 15.21 Only $\text{E}_{\text{effect}}$ 30.70 57.14 17.89 27.24 45.58 15.19 Zero Shot v.v. 27.16 39.65 14.45 24.73 29.76 12.26 Fine-tune v.v. 31.78 65.60 19.39 27.47 47.74 15.65 CEN (Ours) 31.46 63.51 19.25 27.90 49.87 15.76
Figure 4: (a) UMAP visualization of video features learned from CTN (red) and original (blue) captions on MSR-VTT, showing non-overlapping feature spaces. (b) UMAP visualization of video features learned from cause (black) and effect (orange) parts of CTN captions on MSR-VTT, showing near-complete overlap.
Qualitative Analysis: Qualitative examples, in Figure 5, demonstrate CEN’s strength in accurately articulating causal relationships and event sequences, while SOTA methods struggle. For instance, in (a), CEN captures the fatality move’s causal effect, unlike baselines. In (c), CEN links the kick’s precision to scoring, which others miss. Overall, quantitative and qualitative results showcase CEN’s superior performance in understanding and generating causal-temporal narrative video captions compared to existing SOTA methods. More results are provided in Appendix F.
Figure 5: Qualitative examples across scenarios like video games, paper folding, soccer, and singing. CEN (Ours) captions accurately capture causal narratives and temporal sequences from ground truth, outperforming SOTA video captioning methods.
Ablation Results: To examine the effectiveness of CEN, we compare it with 4 baselines in Table 2:
a) $\text{E}_{\text{JointCE}}$: Instead of using two separate video encoders for cause and effect, we train one CLIP-ViT using the combined cause and effect captions separated by space for the text encoder. The performance drops across all metrics and datasets compared to the CEN architecture, which underscores the benefits of dedicated encoders for capturing cause and effect dynamics.
b) Only Ecause and Only Eeffect: These ablations employ only the cause encoder or only the effect encoder, respectively. While retaining reasonable performance, both variants fall short of the performance of the full model across all metrics and datasets.
c) Zero Shot v.v.: In this zero-shot setting, we evaluate the models trained on one dataset (MSVD or MSR-VTT) against the other dataset, without any fine-tuning on the other dataset. The substantial performance drop highlights the importance of dataset-specific fine-tuning for effective causal-temporal narrative learning.
d) Fine-tune v.v.: Building upon the zero-shot case, this variant involves fine-tuning the models trained on one dataset (MSVD or MSR-VTT) against the other dataset. Notably, this fine-tuning process leads to improved performance on the MSVD dataset, even surpassing the CEN model. This observation demonstrates the potential for transfer learning from larger datasets and the ability to leverage their knowledge effectively on smaller datasets through fine-tuning.
Overall, the ablation study validates the efficacy of our proposed CEN architecture, emphasizing the significance of dedicated encoders for capturing cause and effect relationships independently. The performance drops observed in $\text{E}_{\text{JointCE}}$, Only $\text{E}_{\text{cause}}$, and Only $\text{E}_{\text{effect}}$ results, compared to the CEN model, clearly indicate the necessity of explicitly embedding causal-temporal narrative understanding within the architecture through dedicated cause and effect encoders.
5 Limitations
While our work represents a significant step forward in causal-temporal narrative video captioning, our method assumes a certain level of causal and temporal coherence within the video content. Highly complex or convoluted causal relationships may challenge our approach like existing approaches, necessitating further architectural enhancements for improved generalization. Also, the current work primarily focuses on understanding and generating captions that capture causal narratives and temporal dynamics. Explicit integration of spatial reasoning, multi-agent interactions, and long-term dependencies could further enhance the robustness and applicability of our approach. Despite these limitations, our work opens up new avenues for innovative applications and research directions, further solidifying the importance of causal-temporal narrative understanding in video analysis tasks.
6 Conclusion
In this paper, we introduced for the first time a Causal-Temporal Narrative (CTN) captions benchmark and proposed a novel Cause-Effect Network (CEN) tailored for causal-temporal narrative video captioning. This work finds vast applications in automated video summarization, question- answering, assistive technologies, and surveillance. The CTN captions benchmark provides a comprehensive testbed for evaluating video understanding models’ ability to grasp complex temporal and causal dynamics, which will be released for research purpose on https://narrativebridge.github.io/. And CEN explicitly models cause-effect relationships and temporal dynamics to generate rich, contextually relevant descriptions capturing nuanced causal-temporal narrative in videos, demonstrating significant performance improvement over SOTA methods (Section 4). NarrativeBridge lays the foundation for a paradigm shift, where models comprehend underlying causal-temporal narrative driving events, unlocking new frontiers in contextually aware human-machine interactions with video.
For future work, we aim to integrate CTN caption generation with existing image captioning techniques, to annotate unlabeled videos with causal-temporal narrative labels. A few frames of the video will be labelled using off-the-shelf image captioning methods, and CTN caption generation will exploit the labelled frames to generate one coherent caption for the unlabelled video (see Appendix E). This synergistic approach opens new avenues for comprehensive video understanding and annotation, enabling more robust and accurate video analysis pipelines.
7 Acknowledgment
This research was partly supported by the British Broadcasting Corporation Research and Development (BBC R&D), Engineering and Physical Sciences Research Council (EPSRC) Grant EP/V038087/1 “BBC Prosperity Partnership: Future Personalised Object-Based Media Experiences Delivered at Scale Anywhere”.
Appendix
Appendix A Implementation Details
We first extract frames from the video clip and then encode them using the Cause and Effect Video Encoders. We adopt a sampling strategy where we sample one frame per second, with a maximum of 20 frames. This approach ensures that the sampled frames cover the entire duration of the video, providing a comprehensive representation of the video content while maintaining computational efficiency. For the fine-tuning experiment in Table 2, all the hyperparameters remain the same except the learning rate i.e. 0.0000005. For our experimentation of the Cause and Effect Network (CEN), we use standard splits of the MSVD [7] and MSRVTT [55] datasets. For MSVD, we use 1200 videos for training, 100 for validation, and 670 for testing. For MSRVTT, we use 6513 videos for training, 497 for validation, and 2990 for testing. These splits are commonly used in video captioning research to ensure fair comparison with other methods. All the experiments in Stage 1 and Stage 2 of our CEN are run using A100-80GB and RTX 3090-24GB GPUs respectively. We implement SEM-POS [32], AKGNN [17] and GIT [51] using RTX 3090-24GB, A100-80GB and A100-80GB GPUs respectively.
Appendix B Prompt Design Process
In Section 3.1, we introduce our approach for generating Causal-Temporal Narrative (CTN) captions using a large language model (LLM) and few-shot prompting. The prompt design played a crucial role in guiding the LLM to generate captions that accurately capture the cause-effect relationships and temporal dynamics in the video content. Figure 1 showcases the effectiveness of our few-shot based prompt, which resulted in a coherent and contextually relevant CTN caption for the given video. To further illustrate the importance of prompt design and the benefits of few-shot learning, we conduct additional experiments with zero-shot prompting. In these experiments, we evaluate the performance of the LLM in generating CTN captions without providing any example captions in the prompt.
Zero-Shot Prompting: Zero-shot prompting refers to the process of providing large language models (LLMs) with a task prompt without any accompanying examples or demonstrations, requiring the model to generate relevant responses based solely on its pre-trained knowledge and understanding of the prompt. First, we aim to evaluate the performance of LLMs in generating causal-temporal narrative (CTN) captions for videos without providing any few-shot examples.
To achieve this, we construct 4 distinct prompts, each with varying degrees of quality, and experimented with all of them to assess their effectiveness in a zero-shot setting, starting with the most challenging prompt and progressively improving the quality. Prompt 2 presents an extremely terse and poorly structured set of instructions. It contains minimal context and information, merely instructing the model to "Make Cause Effect thing from captions" and listing a set of numbered requirements without providing clear explanations. As expected, the resulting captions (Figure 6) are of low quality and often lack coherence and relevance to the video content. Moreover, the LLM generated multiple captions instead of a single coherent caption, and the output format does not adhere to the specified structure (a single JSON object containing the Cause and Effect statements).
Make Cause Effect thing from captions. Follow: 1. Use only caption info 2. Be relevant 3. No new stuff
4. No conclusions 5. Same time 6. Plain English short sentences 7. 15 words max 8. Just JSON.
[Descriptive Captions]:
<descriptive_captions>
[Causal Temporal Narrative]:
Prompt 2: Zero-shot LLM Prompt with minimal context and unexplained requirements for generating cause-effect from captions.
{ "1": {"Cause": "A car crashes", "Effect": "Guys play beer pong"},
"2": {"Cause": "A car drives through an open field", "Effect": "Kicks up dirt"},
"3": {"Cause": "Unclear", "Effect": "A car flips over"},
"4": {"Cause": "Unclear", "Effect": "A car gets wracked"},
"5": {"Cause": "Unclear", "Effect": "A car is being flipped over"},
"6": {"Cause": "A dirt vehicle rides", "Effect": "It rolls"},
"7": {"Cause": "Unclear", "Effect": "A dune buggy flips over"},
"8": {"Cause": "Unclear", "Effect": "A four wheeler wrecks"},
"9": {"Cause": "A monster truck flips on its side", "Effect": "Several young men shout while playing beer pong"},
"10": {"Cause": "A person drives an offroad car around a field", "Effect": "Unclear"},
"11": {"Cause": "A person flips a go kart", "Effect": "A crowd cheers"},
"12": {"Cause": "Unclear", "Effect": "A race truck is crashing"},
"13": {"Cause": "A truck rolls over itself", "Effect": "Boys cheer on a friend"},
"14": {"Cause": "Unclear", "Effect": "A truck tumbles over on itself"},
"15": {"Cause": "A tumbler crashes on a dirt road", "Effect": "Then a group of guys play beer pong"},
"16": {"Cause": "A type of monster truck crashes", "Effect": "Men are shown celebrating"},
"17": {"Cause": "Unclear", "Effect": "A vehicle flips over"},
"18": {"Cause": "Unclear", "Effect": "An off road vehicle crashes"},
"19": {"Cause": "Unclear", "Effect": "A car crashes while driving"},
"20": {"Cause": "Footage from a monster truck style event", "Effect": "Followed by a frat party"} }
Figure 6: Caption generated using LLM for the example in Figure 1 using Prompt 2
We then progressively improve the prompt quality, moving to a more concise set of instructions in Prompt 3 and a relatively well-structured prompt in Prompt  4. While the captions generated by these improved prompts (Figures 7 and 8) show some enhancements, they still fall short in terms of capturing the causal-temporal narrative accurately and consistently. Additionally, the LLM continues to generate multiple captions and deviate from the desired output format.
You need to create a concise Cause and Effect scenario in JSON format based on some descriptive
captions about a video. The scenario must follow these rules: stick to the given information, be
relevant, dont add new events/details or draw conclusions, maintain time consistency, use
simple language in short sentences, and limit each statement to 15 words. Dont put any extra text
outside the JSON.
[Descriptive Captions]:
<descriptive_captions>
[Causal Temporal Narrative]:
Prompt 3: Zero-shot LLM Prompt with reasonable structure conveying core task of creating cause-effect scenario from captions.
{{"Event": "Car crashes", "Effect": "Guys play beer pong"},
{"Event": "Car drives in open field", "Effect": "Dirt is kicked up"},
{"Event": "Car flips over", "Effect": "None specified"},
{"Event": "Car gets wracked", "Effect": "None specified"},
{"Event": "Car is flipped over", "Effect": "None specified"},
{"Event": "Dirt vehicle rides and rolls", "Effect": "None specified"},
{"Event": "Dune buggy flips over", "Effect": "None specified"},
{"Event": "Four wheeler wrecks", "Effect": "None specified"},
{"Event": "Monster truck flips", "Effect": "Young men shout and play beer pong"},
{"Event": "Person drives offroad car", "Effect": "None specified"},
{"Event": "Person flips go kart", "Effect": "Crowd cheers"},
{"Event": "Race truck crashes", "Effect": "None specified"},
{"Event": "Truck rolls over", "Effect": "Boys cheer on a friend"},
{"Event": "Tumbler crashes on dirt road", "Effect": "Guys play beer pong"},
{"Event": "Monster truck crashes", "Effect": "Men are shown celebrating"},
{"Event": "Vehicle flips over", "Effect": "None specified"},
{"Event": "Off road vehicle crashes", "Effect": "None specified"},
{"Event": "Car crashes while driving", "Effect": "None specified"},
{"Event": "Monster truck event footage", "Effect": "Frat party is shown"}}
Figure 7: Caption generated using LLM for the example in Figure 1 using Prompt 3
You are tasked with generating a causal narrative caption for a video based on a series of provided
descriptive captions. Your generated Cause and Effect scenario should be grounded in the given
information, relevant, avoid introducing new details or implying conclusions, maintain temporal
consistency, use plain English indirect sentences, and limit each Cause and Effect statement to
15 words maximum. Do not include any additional text outside the JSON object.
[Descriptive Captions]:
<descriptive_captions>
[Causal Temporal Narrative]:
Prompt 4: Zero-shot LLM Prompt clearly describing caption-based cause-effect generation task with specified requirements.
{{"Event": "Car crashes", "Effect": "Guys play beer pong"},
{"Event": "Car drives in open field", "Effect": "Dirt is kicked up"},
{"Event": "Car flips over", "Effect": "None specified"},
{"Event": "Car gets wracked", "Effect": "None specified"},
{"Event": "Car is flipped over", "Effect": "None specified"},
{"Event": "Dirt vehicle rides and rolls", "Effect": "None specified"},
{"Event": "Dune buggy flips over", "Effect": "None specified"},
{"Event": "Four wheeler wrecks", "Effect": "None specified"},
{"Event": "Monster truck flips", "Effect": "Young men shout and play beer pong"},
{"Event": "Person drives offroad car", "Effect": "None specified"},
{"Event": "Person flips go kart", "Effect": "Crowd cheers"},
{"Event": "Race truck crashes", "Effect": "None specified"},
{"Event": "Truck rolls over", "Effect": "Boys cheer on a friend"},
{"Event": "Tumbler crashes on dirt road", "Effect": "Guys play beer pong"},
{"Event": "Monster truck crashes", "Effect": "Men are shown celebrating"},
{"Event": "Vehicle flips over", "Effect": "None specified"},
{"Event": "Off road vehicle crashes", "Effect": "None specified"},
{"Event": "Car crashes while driving", "Effect": "None specified"},
{"Event": "Monster truck event footage", "Effect": "Frat party is shown"}}
Figure 8: Caption generated using LLM for the example in Figure 1 using Prompt 4
Finally, we experiment with a zero-shot prompt (Prompt 5) that closely resembles our few-shot based Prompt 1 from Section 3.1, but without the example captions. The captions generated by this prompt (Figure 9) demonstrate improved coherence and relevance compared to the previous zero-shot prompts. However, they still lack the level of detail, accuracy, and contextual understanding exhibited by the captions generated using our few-shot based prompt (Prompt 1). Furthermore, the LLM persists in generating multiple captions.
You are an advanced language model tasked with generating causal temporal narrative captions for a
video. However, you cannot directly access the video series of itself. Instead, you will be provided
with a captions that outline the key events and scenes in the video. Your task is to generate a
concise Cause and Effect scenario, based on the information provided in the descriptive captions.
Be careful, your generated Cause and Effect statements should fulfill the following requirements:
1. Your narrative should be grounded in the information provided by the descriptive captions.
2. Cause and Effect scenario is relevant.
3. It should not introduce any new events or details not mentioned.
4. Avoid implying conclusions.
5. Maintain temporal consistency with the provided captions.
6. Use plain English and direct sentences.
7. Cause and Effect statements each limited to a maximum of 15 words.
8. Do not include any additional text before or after the JSON object.
[Descriptive Captions]:
<descriptive_captions>
[Causal Temporal Narrative]:
Prompt 5: Zero-Shot Prompting for Causal Narrative Caption Generation
{ "Cause and Effect Statements": [ "Car crashes in field, causing it to flip over.", "Dune buggy flips, resulting in a wreck.", "Go kart flipped by person, leading to crowd cheers.", "Race truck crashes, boys cheer in response.", "Truck tumbles, friends shout while playing beer pong.", "Monster truck crashes, men celebrate afterwards.", "Off-road vehicle crashes, causing a tumble.", "Car crash occurs, followed by a game of beer pong." ] }
Figure 9: Caption generated using LLM for the example in Figure 1 using Prompt 5
These experiments highlight the significance of effective prompt design and the advantages of few-shot learning in generating high-quality CTN captions. By providing carefully crafted example captions in the prompt, we enable the LLM to better understand the desired format, style, and level of detail for the generated captions. The few-shot based prompt allows the LLM to learn from and mimic the provided examples, resulting in captions that accurately capture the causal-temporal narrative of the video content.
Crucially, the few-shot based prompt also guides the LLM to generate a single coherent caption in the specified JSON format, ensuring consistency and usability of the generated captions. In contrast, the zero-shot prompts, even when well-structured, rely solely on the LLM’s pre-existing knowledge and understanding of the task description. While the LLM can generate relevant captions to some extent, the lack of explicit examples limits its ability to produce captions with the same level of coherence, accuracy, and contextual understanding as those generated using the few-shot based prompt. Moreover, the zero-shot prompts fail to guide the LLM toward generating a single caption in the desired JSON format, resulting in multiple captions and inconsistent output structures.
These findings underscore the importance of our few-shot based prompt design in the CTN caption generation process and demonstrate its superiority over zero-shot prompting approaches. By leveraging the power of few-shot learning, we enable the LLM to generate high-quality CTN captions that effectively capture the cause-effect relationships and temporal dynamics in video content, as evidenced by the results presented in Figure 1. Additionally, our few-shot based prompt ensures that the generated captions adhere to the specified JSON format, facilitating seamless integration and usability in downstream applications.
Appendix C Comparison of Large Language Models for CTN Caption Generation
To evaluate the effectiveness of large language models (LLMs) in generating high-quality Causal-Temporal Narrative (CTN) captions, we compare the performance of two open source state-of-the-art LLMs at the time of experimentation: Mixtral of Experts [20], which we utilized in our CTN caption generation pipeline, and Llama2-70b [46].
Figure 10: Comparison of Mixtral of Experts and Llama2-70b on a video game sequence involving a fatality move.
We provide the original video captions as input to both models and assess the quality of their generated CTN captions. In the video game example (Figure 10), Mixtral of Experts accurately captures the causal relationship between the fatality move performed by one character and the consequent death of another character in the game. In contrast, Llama2-70b focuses on a specific effect (blood spurting from the character’s neck) without explicitly linking it to the cause.
Figure 11: Comparison of Mixtral of Experts and Llama2-70b on a paper folding tutorial to create a paper airplane.
For the paper folding tutorial (Figure 11), both models correctly identify the causal relationship between folding the paper and creating a paper craft. However, Mixtral of Experts provides a more precise description, specifying that the paper craft being created is a paper airplane.
Figure 12: Comparison of Mixtral of Experts and Llama2-70b on highlights from a soccer match.
In the soccer highlight scenario (Figure 12), Mixtral of Experts successfully captures the causal link between the soccer player’s precise kick and the ball successfully going into the goal. Llama2-70b, on the other hand, mentions the effect (people cheering) without explicitly connecting it to the cause (the player scoring a goal).
Figure 13: Comparison of Mixtral of Experts and Llama2-70b on a boy singing on stage in front of an audience.
Lastly, for the singing performance (Figure 13), both models accurately describe the cause-effect relationship between the boy’s decision to perform on stage and the audience’s reaction of watching and listening. However, Mixtral of Experts provides a more detailed description of the audience’s response.
Overall, the CTN captions generated by Mixtral of Experts consistently demonstrate a better understanding of the causal-temporal narrative in the videos compared to Llama2-70b. Our approach using Mixtral of Experts effectively captures the cause-effect relationships and temporal dynamics, resulting in more accurate and contextually relevant captions.
These results highlight the importance of selecting an appropriate LLM and designing effective prompts for generating high-quality CTN captions. The superior performance of Mixtral of Experts can be attributed to its architecture and training, which enable it to better understand and articulate the complex causal-temporal narratives present in video content.
Appendix D Impact of Automatic Evaluation on CTN Caption Generation
In Section 3.1, we discuss the importance of using automatic evaluation to ensure the quality and relevance of the generated Causal Temporal Narrative (CTN) captions. We employ the EMScore metric to measure the consistency between the generated captions and the video content, discarding captions that fell below a predefined threshold. This appendix visually demonstrates the impact of the automatic evaluation step on the quality of the generated CTN captions.
Figure 14: CTN caption comparison for a video of a female sportscaster interviewing a male athlete.
Figure 15: CTN caption comparison for a video of protests in Brazil due to increased bus ticket prices.
Figure 16: CTN caption comparison for a video discussing a terror attack in Texas.
Figure 17: CTN caption comparison for a video about the Air Force using a surveillance system in Iraq.
Figure 14 shows an example where the CTN caption generated with automatic evaluation accurately captures the causal relationship between the female sportscaster interviewing the male athlete and her asking him questions. In contrast, the caption generated without automatic evaluation is irrelevant to the video content, discussing a man neglecting his allergy medication and suffering a sneezing fit.
Similarly, in Figure 15, the CTN caption generated with automatic evaluation correctly identifies the cause of the protests in Brazil as the increase in bus ticket prices and links it to the effect of a woman being interviewed about her participation in the protests. The caption generated without automatic evaluation, while mentioning the protest and its cause, fails to capture the specific effect of the woman’s interview.
Figure 16 demonstrates how the CTN caption generated with automatic evaluation accurately describes the cause of a terror attack in Texas and its effect on news anchors discussing freedom of speech issues. The caption generated without automatic evaluation is completely unrelated, mentioning a band’s performance and the audience’s reaction.
Lastly, in Figure 17, the CTN caption generated with automatic evaluation correctly links the cause of the Air Force using a surveillance system in Iraq to the effect of a woman talking about the man who created it. The caption generated without automatic evaluation is again irrelevant, discussing a group starting to dance at an event and the atmosphere becoming more lively. These examples clearly illustrate the importance of incorporating automatic evaluation in the CTN caption generation process. By ensuring that the generated captions are consistent with the video content, we can significantly improve the quality and relevance of the CTN captions, enabling the CEN model to learn more meaningful causal and temporal relationships from the videos.
Appendix E Application of CTN Caption Generation for Labeling Unlabeled Videos
Figure 18: Application of our CTN caption generation approach for labeling unlabeled videos. Given an unlabeled video, we extract frames and generate image captions using a state-of-the-art image captioning model (GIT [51]). These captions are then used as input to our LLM-based CTN caption generation pipeline, which produces a CTN caption for the entire video. The generated CTN caption captures the cause-effect relationships and temporal dynamics in the video, enabling effective labeling of unlabeled video content.
Our CTN caption generation approach, as described in Section 3.1, can be effectively applied to the task of labeling unlabeled videos. This application leverages the power of our few-shot based prompt and the LLM’s ability to generate coherent and contextually relevant captions that capture the causal-temporal narrative in video content.
Figure 18 illustrates the pipeline for labeling an unlabeled video using our CTN caption generation approach. First, we extract frames from the unlabeled video. In this example, we extract five frames which are equally spaced in the video. Next, we generate image captions for each of the extracted frames using a state-of-the-art image captioning model. For this demonstration, we employ the GIT [51], which has shown impressive performance in generating accurate captions for individual images. The GIT model generates captions such as "a tractor is driving", "there are four trucks", "a tractor is upside down", "a game is being played", and "two people are celebrating" for the five frames. These image captions serve as the input to our LLM-based CTN caption generation pipeline (see Figure 2).
We use Prompt 1 in Section 3.1, replacing the original descriptive captions with the image captions generated by the GIT model. The LLM then generates a CTN caption based on these image captions, following the specified requirements and format. In this example, the generated CTN caption is: "Cause: ’a tractor driving recklessly flips over and ends up upside down’ Effect: ’causing an accident, while others celebrated unaware’". This caption effectively captures the key events and their causal-temporal relationships in the video, providing a concise and informative summary of the video content.
To ensure the quality and relevance of the generated CTN caption, we employ the same evaluation framework described in Section 3.1. The caption is compared against the video content using the EMScore [41] metric, and only captions that meet a specified threshold are retained. This application demonstrates the versatility and effectiveness of our CTN caption generation approach in labeling unlabeled videos. By leveraging the power of state-of-the-art image captioning models and our few-shot based prompt, we can generate high-quality CTN captions that accurately capture the causal-temporal narrative in video content, even in the absence of human-annotated captions. This approach has the potential to significantly streamline the process of labeling large-scale video datasets and enable more effective video understanding and retrieval tasks.
Appendix F Additional Qualitative Results for CEN Architecture
In this section, we present additional qualitative results comparing the ground truth Causal-Temporal Narrative (CTN) captions with the captions generated by our CEN architecture. These examples further demonstrate the effectiveness of our approach in capturing the causal-temporal relationships and generating accurate and contextually relevant captions.
Figure 19: Comparison of CTN and CEN captions for a video of a person solving a Rubik’s cube.
Figure 20: Comparison of CTN and CEN captions for a video of a man playing the piano.
Figure 21: Comparison of CTN and CEN captions for a video of a man giving a lecture on stage.
Figure 22: Comparison of CTN and CEN captions for a video of a band performing a song.
In Figure 19, both the CTN and CEN captions accurately capture the causal relationship between the person working on the Rubik’s cube and the cube being solved piece by piece. Similarly, in Figure 20, both captions correctly describe the cause-effect relationship between the man playing the piano and the music filling the room.
Figure 21 demonstrates the ability of our CEN architecture to generate captions that capture the audience’s engagement in response to the man’s lecture on stage. While the CTN caption specifically mentions the emotional engagement of the audience due to the sad story, the CEN caption more generally describes the audience listening intently to the lecture.
Lastly, in Figure 22, both the CTN and CEN captions accurately depict the causal relationship between the band performing a song and the crowd enjoying the music. The CEN caption, in particular, directly states the crowd’s enjoyment as a result of the band’s performance.
These additional qualitative examples further validate the effectiveness of our CEN architecture in understanding and articulating the causal-temporal narratives present in videos, generating captions that are coherent, accurate, and contextually relevant.